Give me clearly solution ! Thank you
Three moles of a monatomic gas go from TI = 540K b
Answer #1

A)
del U = f/2 n R del(T)

del U = 3/2 x 3 x 8.314 x 190

del U = 7108.47 J

B)
Q = dU + W

W = Q - dU

W1 = 55000 - 7108.47

W1 = 47891.53 J

W2 = 15000 - 7108.47

W2 = 7891.53 J